# Diffusion Model Patching via Mixture-of-Prompts

### v1

@@ Abstract
We present Diffusion Model Patching (DMP), a simple method to boost the performance of pre-trained diffusion models that have already reached convergence, with a negligible increase in parameters. DMP inserts a small, learnable set of prompts into the model’s input space while keeping the original model frozen. The effectiveness of DMP is not merely due to the addition of parameters but stems from its dynamic gating mechanism, which selects and combines a subset of learnable prompts at every step of the generative process (i.e., reverse denoising steps). This strategy, which we term “mixture-of-prompts”, enables the model to draw on the distinct expertise of each prompt, essentially “patching” the model’s functionality at every step with minimal yet specialized parameters. Uniquely, DMP enhances the model by further training on the same dataset on which it was originally trained, even in a scenario where significant improvements are typically not expected due to model convergence. Experiments show that DMP significantly enhances the converged FID of DiT-L/2 on FFHQ 256$\times$256 by 10.38%, achieved with only a 1.43% parameter increase and 50K additional training iterations.
1 Introduction
@@
The rapid progress in generative modeling has been largely driven by the development and advancement of diffusion models [54, 17], which have garnered considerable attention thanks to their desirable properties, such as stable training, smooth model scaling, and good mode coverage [40]. Diffusion models have set new standards in generating high-quality, diverse samples that closely match the distribution of various datasets [8, 47, 51, 45].
Figure 1: Overview of DMP. We take inspiration from prompt tuning [29] and aim to enhance already converged diffusion models. Our approach incorporates a pool of prompts within the input space, with each prompt learned to excel at certain stages of the denoising process. This is similar to giving a skilled artist an expanded color palette to refine different aspects of their artwork. At every step, a unique blend of prompts (i.e., mixture-of-prompts) is constructed via dynamic gating based on the current noise level. This mechanism is akin to an artist choosing the appropriate color combinations for specific moments. Importantly, our method keeps the diffusion model itself unchanged, and we only use the original training dataset for further training.
@@
Diffusion models are characterized by their multi-step denoising process, which progressively refines random noise into structured outputs, such as images. Each step aims to denoise a noised input, gradually converting completely random noise into a meaningful image. Despite all denoising steps share the same goal of generating high-quality images, each step has distinct characteristics that contribute to shaping the final output [12, 42]. The visual concepts that diffusion models learn vary based on the noise ratio of input [6]. At higher noise levels (timestep $t$ is close to $T$), where images are highly corrupted and thus contents are unrecognizable, the models focus on recovering global structures and colors. As the noise level decreases and images become less corrupted (timestep $t$ is close to 00), the task of recovering images becomes more straightforward, and diffusion models learn to recover fine-grained details. Recent studies [2, 6, 12, 42] suggest that considering stage-specificity is beneficial, as it aligns better with the nuanced requirements of different stages in the generation process. However, many existing diffusion models do not explicitly consider this aspect.
@@
Our goal is to enhance already converged pre-trained diffusion models by introducing stage-specific capabilities. We propose Diffusion Model Patching (DMP), a method that equips pre-trained diffusion models with an enhanced toolkit, enabling them to navigate the generation process with greater finesse and precision. An overview of DMP is shown in Fig. 1. DMP consists of two main components: (1) A small pool of learnable prompts [29], each optimized for particular stages of the generation process. These prompts are attached to the model’s input space and act as “experts” for certain denoising steps (or noise levels). This design enables the model to be directed towards specific behaviors for each stage without retraining the entire model, instead focusing on adjusting only small parameters at the input space. (2) A dynamic gating mechanism that adaptively combines “expert” prompts (or mixture-of-prompts) based on the noise levels of the input image. This dynamic utilization of prompts empowers the model with flexibility, enabling it to utilize distinct aspects of prompt knowledge sets at different stages of generation. By leveraging specialized knowledge embedded in each prompt, the model can adapt to stage-specific requirements throughout the multi-step generation process.
@@
By incorporating these components, we continue training the converged diffusion models using the original dataset on which they were pre-trained. Given that the model has already converged, it is generally assumed that conventional fine-tuning would not lead to significant improvements or may even cause overfitting. However, DMP provides the model with a nuanced understanding of each denoising step, leading to enhanced performance, even when trained on the same data distribution. As shown in Fig. 2, DMP boosts the performance of DiT-L/2 [43] by 10.38% with only 50K iterations on the FFHQ [25] 256 $\times$ 256 dataset.
Figure 2: Further training of fully converged DiT-L/2. DMP achieves a 10.38% gain in FID-50K with only 50K iterations, while others overfit.
@@
Data: DMP boosts model performance using the original dataset, without requiring any external datasets. This is particularly noteworthy as further training of already converged diffusion models on the same dataset typically does not lead to performance gains. DMP differs from general fine-tuning [41, 7], which often transfers knowledge across different datasets.
@@
Computation: DMP patches pre-trained diffusion models by slightly modifying their input space, without updating the model itself. DMP contrasts with methods that train diffusion models from scratch for denoising-stage-specificity [6, 14, 12, 42], which can be computationally expensive and storage-intensive.
@@
Parameter: DMP adds only a negligible number of parameters, approximately 1.26% of the total model parameters (based on DiT-XL/2). This ensures that performance enhancements are achieved cost-effectively.
@@
Model: DMP eliminates the need to train multiple expert networks for different denoising stages. Instead, it combines a few prompts in various ways to learn nuanced behaviors specific to each step. This simplifies the model architecture and training process compared to prior methods [2, 10, 66].
2 Related Work
Diffusion models with stage-specificity.
@@ 2 Related Work
Recent advancements in diffusion models [54, 17, 55] have broadened their utility across various data modalities, including images [47, 50], audios [27], texts [30] and 3D [62], showcasing remarkable versatility in numerous generation tasks. Recent efforts have focused on improving the specificity of denoising stages, with notable progress on both architectural and optimization fronts. (1) On the architectural front, eDiff-I [2] and ERNIE-ViLG 2.0 [10] introduced the concept of utilizing multiple expert denoisers, each tailored to specific noise levels, thereby augmenting the model’s overall capacity. Further, RAPHAEL [66] employed space-time experts, where space experts focus on specific image regions and time experts specialize in particular denoising stages. DTR [42] refined diffusion model architectures by allocating different channel combinations for each denoising step. (2) From an optimization perspective, Choi et al. [6] proposed a strategy to assign higher weights for stages constructing content and lower weights for stages cleaning up the remaining noise. Hang et al. [14] accelerated convergence by framing diffusion training as a multi-task learning problem [4], where loss weights are adjusted based on task difficulty at each timestep. Go et al. [12] mitigated learning conflicts of multiple denoising stages by clustering similar stages based on their signal-to-noise ratios (SNRs). Previous studies aim to improve the specificity of denoising stages, often by assuming either training from scratch or using multiple expert networks, which can be resource-intensive and require significant parameter storage. Our approach achieves stage-specificity without modifying the original model parameters, starting from and using only a single pre-trained diffusion model.
Parameter-efficient Fine-tuning (PEFT) in Diffusion models.
@@
PEFT offers a way to enhance models by tuning a small number of (extra) parameters, avoiding the need to retrain the entire model and significantly reducing computational and storage costs. This is particularly appealing given the complexity and parameter-dense nature of diffusion models [48, 43]. Directly training diffusion models from scratch is often impractical, leading to a growing interest in PEFT methods [63]. Recent advancements in this field can be broadly categorized into three streams: (1) T2i-Adapter [39], SCEdit [23], and ControlNet [69] utilize adapters [19, 20, 5] or side-tuning [68, 57] to modify the neural network’s behavior at specific layers, thereby rewriting activations. (2) Prompt2prompt [15] and Textual Inversion [11] use a concept similar to prompt tuning [31, 29, 37, 22, 71] that modifies input or textual representations to influence subsequent processing without changing the function itself. (3) CustomDiffusion [28], SVDiff [13], and DiffFit [64] focus on partial parameter tuning [67, 65, 32], fine-tuning specific parameters of the neural network, such as bias terms. These approaches have been successful in tuning diffusion models for personalization or customized editing, often using samples different from the original pre-training dataset [15, 49, 28, 13]. These have also been effective in enabling diverse conditional inputs for controllable image generation [39, 23, 69]. In contrast, our work aims to optimize the performance of pre-trained diffusion models with their original training datasets. While being parameter-efficient, our approach targets in-domain enhancements. [^1] Additional related work is in Appendix A.
3 Diffusion Model Patching (DMP)
@@
We propose DMP [^2] Preliminary of diffusion models is in Appendix B., a simple yet effective method aimed at enhancing already converged pre-trained diffusion models by enabling them to leverage knowledge specific to different denoising stages. DMP comprises two key components: (1) a pool of learnable prompts and (2) a dynamic gating mechanism. First, a small number of learnable parameters, referred to as prompts, are attached to the input space of the diffusion model. Second, the dynamic gating mechanism selects the optimal set of prompts (or mixture-of-prompts) based on the noise levels of the input image. Upon these components, we further train the model using the same pre-training dataset to learn prompts while keeping the backbone parameters frozen. We build upon the DiT [43] as our base model. The overall framework of DMP is shown in Fig. 3.
Prompt tuning.
$\hat{{\bm{y}}}^{\prime}={\color[rgb]{0.83203125,0.359375,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.83203125,0.359375,0.26171875}{{\bm{f}}}_{\bm{\theta}^{\prime}}}({\color[rgb]{0.12109375,0.46875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46875,0.70703125}{\bm{x}}}),$ (1)
where $\hat{{\bm{y}}}^{\prime}$ is the refined prediction, ${{\bm{f}}}_{\bm{\theta}^{\prime}}$ is the modified model, and the colors $\bullet$ and $\bullet$ indicate learnable and frozen parameters, respectively. This process is often computationally expensive and resource-intensive, as it requires storing and updating the full model parameters. In contrast, prompt tuning aims to enhance the output $\hat{{\bm{y}}}$ by directly modifying the input ${\bm{x}}$:
$\hat{{\bm{y}}}^{\prime}={\color[rgb]{0.12109375,0.46875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46875,0.70703125}{{\bm{f}}}_{\bm{\theta}}}({\color[rgb]{0.83203125,0.359375,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.83203125,0.359375,0.26171875}{\bm{x}}^{\prime}}).$ (2)
Previous works [31, 29, 22, 71, 61] commonly define ${{\color[rgb]{0.83203125,0.359375,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.83203125,0.359375,0.26171875}{\bm{x}}^{\prime}}}=[{{\color[rgb]{0.83203125,0.359375,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.83203125,0.359375,0.26171875}{\bm{p}}}};{{\color[rgb]{0.12109375,0.46875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46875,0.70703125}{\bm{x}}}}]$, where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation. However, we take a different approach by directly adding prompts to the input, aiming to more explicitly influence the input itself, thus ${{\color[rgb]{0.83203125,0.359375,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.83203125,0.359375,0.26171875}{\bm{x}}^{\prime}}}={{\color[rgb]{0.83203125,0.359375,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.83203125,0.359375,0.26171875}{\bm{p}}}}+{{\color[rgb]{0.12109375,0.46875,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46875,0.70703125}{\bm{x}}}}$. Prompts are optimized via gradient descent, similar to conventional fine-tuning, but without changing the model’s parameters.
Figure 3: DMP framework with DiT [43]. DMP reuses the original dataset used to pre-train DiT.
Motivation.
@@
During the multi-step denoising process, the difficulty and purpose of each stage vary depending on the noise level [6, 14, 12, 42]. Prompt tuning [31, 29, 22] assumes that if a pre-trained model already has sufficient knowledge, carefully constructed prompts can extract knowledge for a specific downstream task from the frozen model. Likewise, we hypothesize that a pre-trained diffusion model already holds general knowledge about all denoising stages, and by learning different mixture-of-prompts for each stage, we can patch the model with stage-specific knowledge.
3.1 Architecture
As our base architecture, we employ DiT [43], which is a transformer-based [59] DDPM [17] operating in the latent space [48]. We use a pre-trained VAE [26] from Stable Diffusion [48] to process input images into a latent code of shape $H\times W\times D$ (for $256\times 256\times 3$ images, the latent code is $32\times 32\times 4$). We then apply noise to latent code. The noisy latent code is divided into $N$ fixed-size patches, each of shape $K\times K\times D$. These patches are linearly embedded, and standard positional encoding [59] is applied to them, resulting in ${\bm{x}}^{(0)}\in\mathbb{R}^{N\times D}$, The patches are then processed by a sequence of $L$ DiT blocks. These blocks are trained with a timestep embedding ${\bm{t}}$ and optionally with class or text embedding. After the last DiT block, the noisy latent patches undergo the final Layer Normalization and are linearly decoded into a $K\times K\times 2D$ tensor ($D$ for noise prediction and another $D$ for diagonal covariance prediction). Finally, the decoded tokens are rearranged to match the original shape $H\times W\times D$.
Learnable prompts.
@@ 3.1 Architecture
Our goal is to efficiently enhance the model with denoising-stage-specific knowledge, focusing on adjusting small parameters within the input space. To achieve this, we start with a pre-trained DiT model and insert $N$ learnable continuous embeddings of dimension $D$, i.e., prompts, into the input space of each DiT block. The set of learnable prompts is denoted as:
@@
Here, ${\bm{p}}^{(i)}$ denotes the prompts for the $i$-th DiT block and $L$ is the total number of DiT blocks in the model. Unlike previous methods [31, 29, 22, 61], where prompts are typically prepended to the input sequence, we directly add them to the input. This offers the advantage of not increasing the sequence length, thereby maintaining nearly the same computation speed as before. Moreover, during the generation process, each prompt added to the input patch provides a direct signal to help denoise specific spatial parts at each timestep. This design choice allows the model to focus on different aspects of the input, aiding in specialized denoising steps. The output of $i$-th DiT block is computed as follows:
[…]
During further training, only the prompts are updated, while the parameters of the DiT blocks remain unchanged.
Dynamic gating.
In Eq. 4, the same prompts are used throughout the training, thus they will learn knowledge that is agnostic to denoising stages. To patch the model with stage-specific knowledge, we introduce dynamic gating. This mechanism blends prompts in varying proportions based on the noise level of an input image. Specifically, we use a timestep embedding ${\bm{t}}$ to represent the noise level at each step of the generation process. For a given ${\bm{t}}$, the gating network $\mathcal{G}$ creates mixtures-of-prompts, thereby redefining Eq. 4 as:
[…]
@@
3.2 Training
Zero-initialization.
We empirically found that random initialization of prompts can disrupt the original information flow early in training, leading to instability and divergence. To ensure stable further training of a pre-trained diffusion model, we start by zero-initializing the prompts. With the prompt addition strategy that we selected before, zero-initialization helps prevent harmful noise from affecting the deep features of neural network layers and preserves the original knowledge at the beginning of training. In the first training step, since prompt parameters are initialized to zero, ${\bm{p}}^{(i-1)}$ term in Eq. 5 evaluates to zero, thus
[…]
As training progresses, the model can gradually incorporate additional signals from the prompts.
Prompt balancing loss.
We adopt two soft constraints from Shazeer et al. [53] to balance the activation of mixtures-of-prompts. (1) Importance balancing: In a mixture-of-experts setup [21, 24], Eigen et al. [9] noted that once experts are selected, they tend to be consistently chosen. In our setup, the importance balancing loss prevents the gating network $\mathcal{G}$ from overly favoring a few prompts with larger weights and encourages all prompts to have similar overall importance. (2) Load balancing: Despite having similar overall importance, prompts might still be activated with imbalanced weights. For instance, one prompt might be assigned with larger weights for a few denoising steps, while another might have smaller weights for many steps. The load balancing loss ensures that prompts are activated almost uniformly across all denoising steps (or noise levels). [^3] For further details about the prompt balancing loss, please refer to Appendix C.
@@ 3.2 Training
DiT-XL/2 675M
+ DMP 683.5M (+1.26%)
Table 1: Parameters. Default image size is 256$\times$256.
Prompt efficiency.
Table 1 presents the model parameters for various versions of the DiT architecture [43] with and without the DMP, ranging from DiT-B/2 to DiT-XL/2 (where “2” denotes the patch size $K$). Assuming a fixed 256$\times$256 resolution, using the DMP increases DiT-B/2 parameters by 1.96%. For the largest model, DiT-XL/2, the use of DMP results in a 1.26% increase to 683.5M parameters. The proportion of DMP parameters to total model parameters decreases as the model size increases, allowing for tuning only a small number of parameters compared to the entire model.
4 Experiments
We evaluate the effectiveness of DMP on various image generation tasks using pre-trained diffusion models that have already converged. Unlike conventional fine-tuning or prompt tuning, we use the same dataset for further training the model that was used in the pre-training phase. We evaluated image quality using FID [16] score, which measures the distance between feature representations of generated and real images using an Inception-v3 model [58].  [^4] Implementation details can be found in Appendix D.
Datasets & Tasks.
We used three datasets for our experiments: (1) FFHQ [25] (for unconditional image generation) contains 70,000 training images of human faces. (2) ImageNet [7] (for class-conditional image generation) involves 1,281,167 training images from 1,000 different classes. (3) MS-COCO [33] (for text-to-image generation) includes 82,783 training images and 40,504 validation images, each annotated with 5 descriptive captions.
FFHQ [7]

### v3

@@ Abstract
We present Diffusion Model Patching (DMP), a simple method to boost the performance of pre-trained diffusion models that have already reached convergence, with a negligible increase in parameters. DMP inserts a small, learnable set of prompts into the model’s input space while keeping the original model frozen. The effectiveness of DMP is not merely due to the addition of parameters but stems from its dynamic gating mechanism, which selects and combines a subset of learnable prompts at every timestep (i.e., reverse denoising steps). This strategy, which we term “mixture-of-prompts”, enables the model to draw on the distinct expertise of each prompt, essentially “patching” the model’s functionality at every timestep with minimal yet specialized parameters. Uniquely, DMP enhances the model by further training on the original dataset already used for pre-training, even in a scenario where significant improvements are typically not expected due to model convergence. Notably, DMP significantly enhances the FID of converged DiT-L/2 by 10.38% on FFHQ, achieved with only a 1.43% parameter increase and 50K additional training iterations.
Introduction
@@
The rapid progress in generative modeling has been largely driven by the development and advancement of diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020), which have garnered considerable attention thanks to their desirable properties, such as stable training, smooth model scaling, and good mode coverage (Nichol and Dhariwal 2021). Diffusion models have set new standards in generating high-quality, diverse samples that closely match the distribution of various datasets (Dhariwal and Nichol 2021; Ramesh et al. 2021; Saharia et al. 2022b; Poole et al. 2022).
@@
Diffusion models are characterized by their multi-step denoising process, which progressively refines random noise into structured outputs, such as images. Each step aims to denoise a noised input, gradually converting completely random noise into a meaningful image. Despite all denoising steps share the same goal of generating high-quality images, each step has distinct characteristics that contribute to shaping the final output (Go et al. 2023; Park et al. 2023). The visual concepts that diffusion models learn vary based on the noise ratio of input (Choi et al. 2022). At higher noise levels (timestep $t$ is close to $T$), where images are highly corrupted and thus contents are unrecognizable, the models focus on recovering global structures and colors. As the noise level decreases and images become less corrupted (timestep $t$ is close to 00), the task of recovering images becomes more straightforward, and diffusion models learn to recover fine-grained details. Recent studies (Balaji et al. 2022; Choi et al. 2022; Go et al. 2023; Park et al. 2023) suggest that considering stage-specificity is beneficial, as it aligns better with the nuanced requirements of different stages in the generation process. However, many existing diffusion models do not explicitly consider this aspect.
Figure 1: Further training of the fully converged DiT-L/2 model using the same dataset as the pre-training phase. Our method, DMP achieves a 10.38% FID improvement in just 50K iterations, while other methods exhibit overfitting.
Figure 2: Overview of DMP. We take inspiration from prompt tuning (Lester, Al-Rfou, and Constant 2021) and aim to enhance already converged diffusion models. Our approach incorporates a pool of prompts within the input space, with each prompt learned to excel at certain stages of the denoising process. At every step, a unique blend of prompts (i.e., mixture-of-prompts) is constructed via dynamic gating based on the current noise level. This mechanism is similar to an skilled artist choosing the appropriate color combinations to refine different aspects of their artwork for specific moments. Importantly, our method keeps the diffusion model itself unchanged, and only use the original training dataset for further training.
@@
Our goal is to enhance already converged pre-trained diffusion models by introducing stage-specific capabilities. We propose Diffusion Model Patching (DMP), a method that equips pre-trained diffusion models with an enhanced toolkit, enabling them to navigate the generation process with greater finesse and precision. An overview of DMP is shown in Fig. 2. DMP consists of two main components: (1) A small pool of learnable prompts (Lester, Al-Rfou, and Constant 2021), each optimized for particular stages of the generation process. These prompts are attached to the model’s input space and act as “experts” for certain denoising steps (or noise levels). This design enables the model to be directed towards specific behaviors for each stage without retraining the entire model, instead adjusting only small parameters at the input space. (2) A dynamic gating mechanism that adaptively generates “expert” prompts (or mixture-of-prompts) based on the noise levels of the input image. This dynamic utilization of prompts empowers the model with flexibility, enabling it to utilize distinct aspects of prompt knowledge sets at different stages of generation. By leveraging specialized knowledge embedded in each prompt, the model can adapt to stage-specific requirements throughout the multi-step generation process.
@@
By incorporating these components, we continue training the converged diffusion models using the original dataset on which they were pre-trained. Given that the model has already converged, it is generally assumed that conventional fine-tuning would not lead to significant improvements or may even cause overfitting. However, DMP provides the model with a nuanced understanding of each denoising step, leading to enhanced performance, even when trained on the same data distribution. As shown in Fig. 1, DMP boosts the performance of DiT-L/2 (Peebles and Xie 2022) by 10.38% with only 50K iterations on the FFHQ (Karras, Laine, and Aila 2019) 256 $\times$ 256 dataset.
@@
Data: DMP boosts model performance using the original dataset, without requiring any external datasets. This is particularly noteworthy as further training of already converged diffusion models on the same dataset typically does not lead to performance gains. DMP differs from general fine-tuning (Pan and Yang 2009; Deng et al. 2009), which often transfers knowledge across different datasets.
@@
Computational Efficiency: DMP patches pre-trained diffusion models by slightly modifying their input space, without updating the model itself. DMP contrasts with methods that train diffusion models from scratch for denoising-stage-specificity (Choi et al. 2022; Hang et al. 2023; Go et al. 2023; Park et al. 2023, 2024), which can be computationally expensive and storage-intensive.
@@
Parameter Efficiency: DMP adds only a negligible number of parameters, approximately 1.43% of the total model parameters (based on DiT-L/2). This ensures that performance enhancements are achieved cost-effectively.
@@
Model: DMP eliminates the need to train multiple expert networks for different denoising stages. Instead, it combines a few prompts in various ways to learn nuanced behaviors specific to each step. This simplifies the model architecture and training process compared to prior methods (Balaji et al. 2022; Feng et al. 2023; Xue et al. 2023; Park et al. 2024).
Related Work
@@
Diffusion models with stage-specificity. Recent advancements in diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) have broadened their utility across various data modalities, including images (Ramesh et al. 2021; Saharia et al. 2022a), audios (Kong et al. 2020), texts (Li et al. 2022) and 3D (Woo et al. 2023), showcasing remarkable versatility in numerous generation tasks. Recent efforts have focused on improving the specificity of denoising stages, with notable progress on both architectural and optimization fronts. (1) On the architectural front, eDiff-I (Balaji et al. 2022), ERNIE-ViLG 2.0 (Feng et al. 2023), and RAPHAEL (Xue et al. 2023) introduced the concept of utilizing multiple expert denoisers, each tailored to specific noise levels, thereby augmenting the model’s capacity. DTR (Park et al. 2023) refined diffusion model architectures by allocating different channel combinations for each denoising step. (2) From an optimization perspective, P2 Weight (Choi et al. 2022) and Min-SNR Weight (Hang et al. 2023) accelerated convergence by framing diffusion training as a multi-task learning problem (Caruana 1997), where loss weights are adjusted based on task difficulty at each timestep. Go et al. (Go et al. 2023) mitigated learning conflicts of multiple denoising stages by clustering similar stages based on their signal-to-noise ratios (SNRs). Previous studies aim to improve the specificity of denoising stages, often by assuming either training from scratch or using multiple expert networks, which can be resource-intensive and require significant parameter storage. Whereas, our approach achieves stage-specificity without modifying the original model parameters, starting from and using only a single pre-trained diffusion model.
@@
Parameter-efficient Fine-tuning (PEFT) in Diffusion models. PEFT offers a way to enhance models by tuning a small number of (extra) parameters, avoiding the need to retrain the entire model and significantly reducing computational and storage costs (Xiang et al. 2023). This is particularly appealing given the complexity and parameter-dense nature of diffusion models (Rombach et al. 2022; Peebles and Xie 2022), where directly training diffusion models from scratch is impractical. Recent advancements in this field can be broadly categorized into three streams: (1) T2i-Adapter (Mou et al. 2023), SCEdit (Jiang et al. 2023), ControlNet (Zhang, Rao, and Agrawala 2023) and CDMs (Golatkar et al. 2023) utilize adapters (Houlsby et al. 2019; Hu et al. 2021; Chen et al. 2022) or side-tuning (Zhang et al. 2020; Sung, Cho, and Bansal 2022) to modify the neural network’s behavior at specific layers. (2) Textual Inversion (Hertz et al. 2022; Gal et al. 2022) use a concept similar to prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021; Logan IV et al. 2021; Jia et al. 2022; Zhou et al. 2022b) that modifies input or textual representations to influence subsequent processing without changing the function itself. (3) CustomDiffusion (Kumari et al. 2023), SVDiff (Han et al. 2023), and DiffFit (Xie et al. 2023) focus on partial parameter tuning (Zaken, Ravfogel, and Goldberg 2021; Xu et al. 2021; Lian et al. 2022), fine-tuning specific parameters of the neural network, such as bias terms. These approaches have been successful in tuning diffusion models for personalization, using samples different from the original pre-training dataset. In contrast, our work aims to optimize the performance of pre-trained diffusion models with their original training datasets. While being parameter-efficient, our approach targets in-domain enhancements. [^1] Extended related work is in Appendix.
Diffusion Model Patching (DMP)
Figure 3: DMP framework with DiT (Peebles and Xie 2022). DMP is designed to adaptively generate optimal prompts tailored to specific timesteps. DMP uses the original training dataset—previously used for pre-training diffusion models—for fine-tuning. Operating entirely through prompt-based tuning in the input space, DMP eliminates the need for modifications to either the model architecture or the overall training process, ensuring seamless integration and efficiency.
@@
We propose DMP, a simple yet effective method aimed at enhancing already converged pre-trained diffusion models by enabling them to leverage knowledge specific to different denoising stages. [^2] See preliminaries in Appendix. DMP comprises two key components: (1) a pool of learnable prompts and (2) a dynamic gating mechanism. First, a small number of learnable prompts are attached to the input space of the diffusion model. Second, the dynamic gating mechanism selects the optimal set of prompts (or mixture-of-prompts) based on the noise levels of the input image. Upon these components, we further train the model using the same pre-training dataset to learn prompts while keeping the backbone parameters frozen. The overall framework of DMP is shown in Fig. 3.
@@
Motivation. During the multi-step denoising process, the difficulty and purpose of each stage vary depending on the noise level (Choi et al. 2022; Hang et al. 2023; Go et al. 2023; Park et al. 2023). Prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021; Jia et al. 2022) assumes that if a pre-trained model already has sufficient knowledge, carefully constructed prompts can extract knowledge for a specific downstream task from the frozen model. Likewise, we hypothesize that a pre-trained diffusion model already holds general knowledge about all denoising stages, and by learning different mixture-of-prompts for each stage, we can patch the model with stage-specific knowledge.
Architecture. As our base architecture, we employ DiT (Peebles and Xie 2022), which is a transformer-based (Vaswani et al. 2017), and Stable Diffusion (Rombach et al. 2022), which is a UNet-based (Ho, Jain, and Abbeel 2020), both operating in the latent space (Rombach et al. 2022). We use a pre-trained VAE (Kingma and Welling 2013) from Stable Diffusion (Rombach et al. 2022) to process input images into a latent code of shape $H\times W\times D$ (for $256\times 256\times 3$ images, the latent code is $32\times 32\times 4$). For DiT, the noisy latent code is divided into $N$ fixed-size patches, each of shape $K\times K\times D$. Our DMP method adaptively adjusts the size of learnable prompts to the input size of each model, where $N\times D$ for DiT and $H\times W\times D$ for Stable Diffusion.
@@
Learnable prompts. Our goal is to efficiently enhance the model with denoising-stage-specific knowledge, adjusting small parameters within the input space. To achieve this, we start with a pre-trained DiT model as an example for explanation. Firstly, we insert $N$ learnable continuous embeddings of dimension $D$, i.e., prompts, into the input space of each DiT block. The set of learnable prompts is denoted as:
@@
Here, ${\bm{p}}^{(i)}$ denotes the prompts for the $i$-th DiT block and $L$ is the total number of DiT blocks in the model. Unlike previous methods (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021; Jia et al. 2022; Wang et al. 2022), where prompts are typically prepended to the input sequence, we directly add them to the input. This offers the advantage of not increasing the sequence length, thereby maintaining nearly the same computation speed as before. Moreover, during the generation process, each prompt added to the input patch provides a direct signal to help denoise specific spatial parts at each timestep. This design choice allows the model to focus on different aspects of the input at each timestep, aiding in specialized denoising steps. The output of $i$-th DiT block is computed as follows:
${\bm{x}}^{(i+1)}={\rm Block}^{(i)}_{frozen}({\bm{p}}^{(i)}_{learn}+{\bm{x}}^{(i)}).$ (2)
During further training, only the prompts are updated, while the DiT blocks remain unchanged. The subscript $learn$ indicates learnable parameters, while $frozen$ indicates frozen parameters.
Dynamic gating. In Eq. 2, the same prompts are used throughout the training, thus they will learn denoising-stage-agnostic knowledge. To patch the model with stage-specific knowledge, we introduce dynamic gating. This mechanism blends prompts in varying proportions based on the noise level of an input image. Specifically, we use a timestep embedding ${\bm{t}}$ to represent the noise level at each step of the generation process. For a given ${\bm{t}}$, the gating network $\mathcal{G}$ creates the stage-specific mask of shape $N\times 1$ used for generating mixtures-of-prompts, thereby redefining Eq. 2 as:
${\bm{x}}^{(i+1)}={\rm Block}^{(i)}_{frozen}\big{(}{\sigma}(\mathcal{G}_{learn}([{\bm{t}};i]))\odot{\bm{p}}^{(i)}_{learn}+{\bm{x}}^{(i)}\big{)},$ (3)
@@
Training
Zero-initialization. We empirically found that random initialization of prompts disrupts the early training process, leading to instability and divergence. To ensure stable further training of a pre-trained diffusion model, we start by zero-initializing the prompts. With the prompt addition strategy that we selected before, zero-initialization prevents harmful noise from affecting the deep features of neural network layers and preserves the original knowledge at the beginning of training. As training progresses, the model gradually incorporate additional signals from the prompts.
Prompt balancing loss. We adopt two soft constraints from Shazeer et al. (2017) to balance the activation of mixtures-of-prompts. (1) Load balancing: In a mixture-of-experts setup (Jacobs et al. 1991; Jordan and Jacobs 1994), Eigen et al. (Eigen, Ranzato, and Sutskever 2013) noted that once experts are selected, they tend to be consistently chosen. In our setup, the load balancing loss prevents the gating network $\mathcal{G}$ from overly favoring a few prompts with larger weights and encourages all prompts to uniformly selected. (2) Importance balancing: Despite having similar overall load, prompts might still be activated with imbalanced weights. For instance, one prompt might be assigned with larger weights for a few denoising steps, while another might have smaller weights for many steps. The load balancing loss ensures that prompts are activated with similar overall importance across all denoising steps. [^3] For further details about prompt balancing loss, see Appendix.
@@
SD v1.5 890M
+ DMP 892M (+0.32%)
Table 1: Parameters. Default image size is 256$\times$256. SD v1.5 indicates Stable Diffusion v1.5 (Rombach et al. 2022).
Prompt efficiency. Table 1 presents the model parameters for various versions of the DiT architecture (Peebles and Xie 2022) ranging from DiT-B/2 to DiT-L/2 (where “2” denotes the patch size $K$) and Stable Diffusion v1.5 with and without the DMP. Assuming a fixed 256$\times$256 resolution, using the DMP increases DiT-B/2 parameters by 1.96%. For the largest model, Stable Diffusion v1.5, the use of DMP results in a 0.3% increase to 892M parameters. The proportion of DMP parameters to total model parameters decreases as the model size increases, allowing for tuning only a small number of parameters compared to the entire model.
Resolution ($256\times 256$) FFHQ